Find equation of line passing through the point: (5,-2) and (-3,4)
Please show me the steps to get the answer. Thanks

Answer by rfer(16322)

slope m=(4--2)/(-3-5)
m=6/-8
m=-3/4
------------
y=mx+b
-2=-3/4*5+b
-2=-15/4+b
-8/4+15/4=b
7/4=b
b=1 3/4
-------------
y=-3/4x+1 3/4